# Curve fitting on a quantum annealer for an advanced navigation method

### v1

@@ III.1 Linear regression
where the ordinate values are distorted by artificial noise. At each $y_{i}$, it is a sample drawn from a normal distribution centered about zero with a standard deviation of $0.02$.
Regarding the expansion $f(x)=\sum_{j=0}^{m-1}c_{j}\phi_{j}(x)$ of the fit function, linear regression corresponds to $m=2$ and $\phi_{0}(x)=1$, $\phi_{1}(x)=x$ such that $f(x)=c_{0}+c_{1}x$ with regression parameters $c_{0}$ and $c_{1}$. For the corresponding QUBO problem, we choose $L=7$ and $U=1$, which is sufficient for the expected order of magnitude of the regression parameters. Since the latter are expected to be positive, the general QUBO problem of size $(2md)\times(2md)$ can be reduced to $(md)\times(md)$, yielding a matrix of size $18\times 18$ for this problem.
In the left diagram of Fig. 1, we show the data as given in Eq. 26 (gray dots) together with the optimal regression lines $f^{\ast}(x)=c_{0}^{\ast}+c_{1}^{\ast}x$ with optimal coefficients $c_{0}^{\ast}$ and $c_{1}^{\ast}$ obtained classically (solid blue), i.e., according to Section II.1 (specifically, using Eq. 9), and by solving the corresponding QUBO formulation via tabu search on a conventional computer (dashed red) as well as through quantum annealing on a D-Wave device (dash-dotted green). [^2] We used D-Wave’s Advantage quantum processing unit. For all numerical experiments performed here, we used the same parameters for the quantum annealing device. Particularly, we performed 32 runs with 128 reads each to collect potential solution candidates and finally took the lowest-energy solution. Comparing the three solutions, we find that the results for the regression lines agree up to the second decimal point. Moreover, the solution obtained via tabu search is also found by quantum annealing, their results are identical. In order to gauge the quality of the quantum annealing solution, in the right diagram of Fig. 1 we show the relative error
$\varepsilon_{\textup{rel}}(a,b)=\frac{\lvert a-b\rvert}{\max\{\lvert a\rvert,\delta\}}\qquad(a,b\in\mathbb{R},\,\delta>0)$ (27)
between the regression lines obtained classically and through quantum annealing. It is below one percent for all abscissae of the data. For the sake of completeness, in Tab. 1 we show the results for the optimal regression parameters.
@@
Thus, in the case of linear regression, our results show that formulating curve fitting as a QUBO problem indeed works and yields a solution on-par with the classical one. They also suggest that quantum annealing in its current state is (within limits) able to deliver promising and practically usable results for this continuous problem.
$c_{0}^{\ast}$ $c_{1}^{\ast}$
Classical solution $0.24377$ $0.50474$
Quantum annealing $0.24219$ $0.50781$
$\varepsilon_{\textup{rel}}$ [%] $0.65$ $0.61$
Table 1: Optimal regression parameters obtained classically and via quantum annealing.
III.2 Piecewise-linear approximation
Figure 2: Left: data (gray dots) together with optimal piecewise-linear approximations obtained classically ($f_{\textup{cl}}^{\ast}$, solid blue) and by a QUBO solved with tabu search ($f_{\textup{tb}}^{\ast}$, dashed red) as well as quantum annealing ($f_{\textup{qa}}^{\ast}$, dash-dotted green). Right: relative errors between the classical solution and the tabu search as well as quantum annealing solutions.
We continue with an example where linear regression is not appropriate. Inspired by observations we have made during the analysis of fuel-consumption data of a container vessel, we now consider the data
$x_{i}=\frac{i}{n-1}\,,\qquad y_{i}=\frac{1}{4}x_{i}+\frac{3}{4}x_{i}^{3}+\frac{1}{10}+\textup{noise}\,,$ (28)
again with $n=32$, and the noise are samples drawn from a normal distribution. These data points shall be approximated by a piecewise linear function. To this end, we choose $m=4$ and use triangular functions for the function expansion, $\phi_{j}=\Lambda_{j}$ (see Eq. 4); their supporting points $\tilde{x}_{0},\dotsc,\tilde{x}_{3}$ are linearly distributed over the abscissa interval $[0,1]$. For the QUBO problem, we choose $L=8$ and $U=1$.
In the left diagram of Fig. 2, we display the optimal fit $f^{\ast}(x)=c_{0}^{\ast}\Lambda_{0}(x)+\dotso+c_{3}^{\ast}\Lambda_{3}(x)$ obtained classically according to Eq. 9 (solid blue) and by solving the QUBO via tabu search (dashed red) as well as quantum annealing (dash-dotted green). Compared to linear regression, the tabu solution is again not distinguishable from the classical one by the eye while the quantum annealing solution shows slight deviations. However, the latter are rather small, and although the QUBO matrix is more than twice as large ($40\times 40$) as for linear regression, the overall fit quality of the quantum solution is still satisfying. This is also reflected in the relative errors between the classical and QUBO solutions that are below ten percent across the whole abscissa range (right diagram of Fig. 2). In Tab. 2, we show the results for the optimal coefficients.
$c_{0}^{\ast}$ $c_{1}^{\ast}$ $c_{2}^{\ast}$ $c_{3}^{\ast}$
Classical $0.09149$ $0.19231$ $0.45782$ $1.06377$
Quantum $0.09375$ $0.20703$ $0.44922$ $1.07031$
$\varepsilon_{\textup{rel}}$ [%] $2.47$ $7.65$ $1.88$ $0.61$
Table 2: Optimal coefficients of the individual triangular functions obtained classically and via quantum annealing.
@@ III.3 Chebyshev polynomials
Figure 3: Left: data (gray dots) together with optimal Chebyshev approximations obtained classically ($f_{\textup{cl}}^{\ast}$, solid blue) and by a QUBO solved with tabu search ($f_{\textup{tb}}^{\ast}$, dashed red) as well as quantum annealing ($f_{\textup{qa}}^{\ast}$, dash-dotted green). Right: relative errors between the classical solution and the tabu search as well as quantum annealing solutions.
@@
Figure 4: Heat map plots of the QUBO matrices $Q$ of the piecewise-linear approximation (left) and Chebyshev polynomial approximation (right) as discussed in Sections III.2 and III.3, respectively. Here, we display the upper-triangularized versions of these matrices since every QUBO problem can be written in a way such that $Q$ is an upper-triangular matrix.
Last, we use Chebyshev polynomials of the first kind for fitting the data described in Eq. 27, $\phi_{j}=T_{j}$ (see Eq. 3), again with $m=4$. Thus, the fit function reads explicitly
$f(x)=c_{0}+c_{1}x+c_{2}\hskip 0.83344pt(2x^{2}-1)+c_{3}\hskip 0.83344pt(4x^{3}-3x)\,.$ (29)
For the corresponding QUBO formulation, we use the same precision as before ($L=8$ and $U=1$).
As apparent from the results shown in Fig. 3, the tabu solution of the QUBO problem is again practically identical to the classical one. On the other hand, the quantum solution now shows visible deviations in form of under- and overshooting with respect to the tabu and classical solutions. The corresponding relative error is about ten percent while the one for the tabu solution is roughly one magnitude smaller.
Thus, even though the QUBO matrix is of the same size as for the piecewise-linear approximation and constructed from the same underlying set of data points, the resulting quantum annealing solution for the Chebyshev approximation is of lower quality than the one for a piecewise-linear curve fitting using triangular functions. We find that this is also the case if we change the precision of the binary decomposition of the expansion coefficients (see Eq. 11) by varying $L$ and/or $U$, i.e., for smaller or larger QUBO matrices, and we were not able to find a quantum annealing solution with the same quality as the ones for linear regression and piecewise-linear approximation. From our point of view though, the quantum result shown in Fig. 3 is still reasonable.
Alas, for a QUBO matrix with a size of about $60\times 60$ and larger, we were not able to find a satisfying solution for the Chebyshev approximation through quantum annealing at all, while a high-quality solution can still be found via tabu search. For the said size of the QUBO problem, finding a satisfying piecewise-linear approximation solely using quantum annealing (i.e., with a relative error of about ten percent compared to the tabu solution) is also rather difficult. From our point of view, this seems to hint toward the limitations of the current quantum annealing hardware for this problem.
@@
More precisely, the QUBO problem, which can be seen as a graph with logical qubits representing nodes that are connected by edges with weights given by the matrix elements of the QUBO matrix, needs to be mapped onto the so-called working graph of the quantum annealing device. Therefore, one might need several physical qubits to represent one logical qubit. Particularly for large and densly-connected QUBO problems, this could yield an Ising model that exhibits a complicated low-energy landscape where the ground state is difficult to access. This would explain why we were not successful in obtaining a feasible Chebyshev approximation as soon as its QUBO matrix is larger than the above mentioned size of about $60\times 60$ because the matrix is, as apparent from the right diagram of Fig. 4, indeed relatively highly connected.
@@
Finally, we would like to emphasize that by solving the QUBO problems for piecewise-linear and Chebyshev approximations using a heuristic technique, e.g., tabu search or simulated annealing, we (almost) always find a solution practically on-par with the classical one—even for sizes of the QUBO matrix for which the quantum annealing solution turned out to be unpractical. Thus, we are confident that a next-generation quantum annealing device will be capable of solving these larger QUBO problems in the future.

### v2

@@ III.1 Sample data
which are all distorted by noise $\varepsilon_{i}$. At each ordinate $y_{i}$, the latter is a random variable drawn from a normal distribution centered about zero with variance $(0.03)^{2}$, i.e., $\varepsilon_{i}\sim\mathcal{N}(\mu=0,\sigma^{2}=(0.03)^{2})$ for $i=0,\dotsc,n-1$.
In order to compare results, we use either the absolute percentage error (APE)
$\operatorname{ape}(a,b)=\frac{\lvert a-b\rvert}{\max\{\lvert a\rvert,\delta\}}\times 100$ (25)
with $a,b\in\mathbb{R}$, $\delta>0$ or the root mean squared error (RMSE)
$\operatorname{rmse}(f^{\ast})=\sqrt{\frac{1}{n}\sum_{i=0}^{n-1}(y_{i}-f^{\ast}(x_{i}))^{2}}\,,$ (26)
where $f^{\ast}(x)=\sum_{j=0}^{m-1}c_{j}^{\ast}\phi_{j}(x)$ is a given optimal fit with optimal expansion coefficients $c_{j}^{\ast}$ obtained from solving the least-squares optimization problem (see Eqs. 1 and 2). Equation 26 aggregates the squared differences between the actual and approximated values for each sample data point—the very same measure that is minimized by the method of least squares—and is thus well-suited to compare optimal least-squares fits $f^{\ast}$ obtained via different techniques (see Sections II.1 and II.2).
Furthermore, we always normalize the ordinate values to the unit interval before the optimal fit is computed. That is, the least-squares method actually operates on min-max-normalized data points $\{(x_{i},y_{i}^{\prime})\}_{i=0}^{n-1}$ with
$y_{i}^{\prime}=\frac{y_{i}-y_{\textup{min}}}{y_{\textup{max}}-y_{\textup{min}}}\in[0,1]\,,$ (27)
where $y_{\textup{min}}=\min_{i}y_{i}$ and $y_{\textup{max}}=\max_{i}y_{i}$. Thus, all optimal expansion coefficients $c_{j}^{\ast}$ presented here are understood to be obtained with respect to min-max-normalized data $y_{i}^{\prime}$.
Now, in the following we consider three different standardized functions $\phi_{j}$ for the expansion 2 of the fit function, apply them to our sample data, and discuss the quality of the solutions obtained via different strategies as well as the influence of certain parameters.
III.2 Triangular functions
We begin with triangular functions as defined in Eq. 4, $\phi_{j}=\Lambda_{j}$, i.e., the fit function is given by
$f(x)=c_{0}\Lambda_{0}(x)+\dotsb+c_{m-1}\Lambda_{m-1}(x)\,,$ (28)
and the least squares optimization problem thus explicitly reads
$\min_{c\in\mathbb{R}^{m}}\sum_{i=0}^{n-1}\,\Biggl{(}y_{i}-\sum_{j=0}^{m-1}c_{j}\Lambda_{j}(x_{i})\Biggr{)}^{2}\,.$ (29)
First, we consider the simple yet important case of linear regression. To this end, we choose $n=64$ and the linear sample data 21. Furthermore, we set $m=2$, which implies that the two supporting points of the triangular functions (see Eq. 4) are simply the first and last abscissa values 20 of the sample data, respectively, i.e., $\tilde{x}_{0}=x_{0}=0$ and $\tilde{x}_{1}=x_{n-1}=1$. This results in a fit function $f(x)=c_{0}\Lambda_{0}(x)+c_{1}\Lambda_{1}(x)$ that is a straight line across the whole abscissa interval $[0,1]$ of our sample data. For the corresponding QUBO formulation of the problem, we choose $d=10$ and $p=d-2=8$, which results in a QUBO matrix of size $md\times md=20\times 20$.
Table 1: Optimal linear regression parameters obtained classically and by solving a QUBO. Also shown are the absolute percentage errors (APE) of the optimal parameters obtained from the QUBO with respect to the classical ones.
$c_{0}^{\ast}$ $c_{1}^{\ast}$ APE $c_{0}^{\ast}$ [%] APE $c_{1}^{\ast}$ [%]
Classical (matrix inversion) $0.010664$ $0.978332$
QUBO (tabu search) $0.010742$ $0.978516$ $0.73$ $0.02$
QUBO (quantum annealing) $0.011719$ $0.977539$ $9.89$ $0.08$
Table 2: Same as Tab. 3 but with the linear sample data now given by $y_{i}=-x_{i}/4+1/3+\varepsilon_{i}$.
$c_{0}^{\ast}$ $c_{1}^{\ast}$ APE $c_{0}^{\ast}$ [%] APE $c_{1}^{\ast}$ [%]
Classical (matrix inversion) $0.874487$ $0.200203$
QUBO (tabu search) $0.875000$ $0.200195$ $0.06$ $0.004$
QUBO (quantum annealing) $0.872070$ $0.196289$ $0.28$ $1.96$
Table 3: APE between the RMSE of the optimal classical solution $f_{\textup{cl}}^{\ast}$ and the optimal QUBO solutions $f_{\textup{tb}}^{\ast}$ and $f_{\textup{qa}}^{\ast}$ obtained via tabu search and quantum annealing, respectively, for different values of $n$. The underlying sample data set is the linear one (Eq. 21) and a linear regression is used for the approximation.
$n$ $\operatorname{ape}(\operatorname{rmse}(f_{\textup{cl}}^{\ast}),\operatorname{rmse}(f_{\textup{tb}}^{\ast}))$ [%] $\operatorname{ape}(\operatorname{rmse}(f_{\textup{cl}}^{\ast}),\operatorname{rmse}(f_{\textup{qa}}^{\ast}))$ [%]
$64$ $2.88\times 10^{-4}$ $2.23\times 10^{-2}$
$128$ $3.64\times 10^{-4}$ $4.50\times 10^{-2}$
$256$ $1.10\times 10^{-4}$ $1.22\times 10^{-2}$
$512$ $1.09\times 10^{-4}$ $1.71\times 10^{-2}$
$1024$ $6.80\times 10^{-4}$ $3.16\times 10^{-2}$
In Tab. 3, we show our results for the optimal linear regression parameters $c_{0}^{\ast}$ and $c_{1}^{\ast}$ obtained classically by (naive) matrix inversion according to Eq. 9 and by solving the corresponding QUBO as given in Eq. 16; we also show the APE (see Eq. 25) of the optimal coefficients obtain via a QUBO with respect to the classical results. The QUBO is solved by means of a tabu search as well as by quantum annealing on a D-Wave quantum processing unit. [^2] For the tabu search, we used the MST2 multi-start tabu algorithm of Ref. [57]. Regarding quantum annealing for solving the QUBO, we used D-Wave’s Advantage2 system. In particular, we performed 100 runs with 100 reads each with an annealing time of 200 microseconds to collect potential solution candidates, yielding 10000 samples in total, and finally took the lowest-energy solution. This setup is used throughout the whole section whenever a QUBO problem is solved via tabu search or quantum annealing. Comparing the three solutions, we find that all agree up to the second decimal point. The QUBO solution found by the tabu search is particularly close to the exact classical solution, agreeing up to the third decimal point. Though the quantum annealing solution differs more noticeable, it is still in good agreement with the two other results. Furthermore, we find qualitatively similar results if we replace the numerical coefficients in the linear sample data 21 by other values, e.g., randomly drawn from $[-1,1]$. As an example, Tab. 3 shows the results for a different set of linear sample data.
@@ III.2 Triangular functions
Thus, in the case of linear regression, our results indicate that formulating curve fitting as a QUBO problem indeed works and yields a solution in satisfying agreement with the classical one. They also suggest that quantum annealing in its current state is in principle able (within limits) to deliver promising and practically usable results for this continuous problem. [^3] Because of the probabilistic (or stochastic) nature of quantum annealing and the fact that the quantum annealing device is a real (i.e., non-ideal) system—thus susceptible to thermal fluctuations and interactions with its environment—subsequent calculations with the same input may yield slightly different results. Based on our experiences and observations in the course of this work, the quantum annealing results presented here seem to be accompanied with an overall error of about five to ten percent.
It is furthermore instructive to investigate the influence of certain parameters on the solutions. In Tab. 3, we show the APE between the RMSE of the optimal classical solution $f_{\textup{cl}}^{\ast}$ and the optimal QUBO solutions $f_{\textup{tb}}^{\ast}$ and $f_{\textup{qa}}^{\ast}$ obtained via tabu search and quantum annealing, respectively, for different values of the number of sample data points $n$ (again for the case of linear regression with sample data 21). We conclude that the number of to-be-approximated data points has a negligible influence on the solution quality. This is expected because $n$ enters only implicitly in the course of determining the entries of the QUBO matrix (see Eqs. 5 and 6) while its size is given by $md\times md$, see Eq. 17, i.e., only determined by $m$ and $d$. We find the same result if we use our other sample data sets instead of the linear one. Therefore, we fix $n=64$ for the remainder of this section.
In Fig. 1, we show the RMSE of the tabu (red squares) and quantum annealing (green diamonds) solutions $f_{\textup{tb}}^{\ast}$ and $f_{\textup{qa}}^{\ast}$, respectively, as a function of $d$, the number of digits that is used for the binary fixed-point format of the coefficients $c_{j}$, see Eq. 11, with $p=d-1$ fixed. Furthermore, we show the RMSE of the classical solution $f_{\textup{cl}}^{\ast}$, which is exact in the sense of numerical precision and thus independent of $d$. Here, we use the cubic sample data 23 and $m=4$, i.e., we are beyond linear regression and have a truly piecewise-linear fit of the form $f(x)=c_{0}\Lambda_{0}(x)+\dotsb+c_{3}\Lambda_{3}(x)$ that is not just a straight regression line. For $d\geq 8$, the tabu solutions are very close to the classical one and a further increase of $d$ has only a negligible impact. The same holds for the quantum annealing solutions, though they are always less accurate than the corresponding tabu results. Both are only identical provided the QUBO matrix is small enough, namely for $d\leq 5$, which corresponds to a QUBO matrix of size $20\times 20$ or smaller. In our opinion, however, the quantum annealing solutions for larger QUBO sizes are still in acceptable agreement with the tabu and classical ones. We find qualitatively similar results if we use the quadratic sample data or if the numerical coefficients of the cubic data are randomly drawn from $[-1,1]$. Therefore, $d=8$ (with $p=d-1$) seems to be sufficient for the binary fixed-point representation of the expansion coefficients $c_{j}$ regarding the cases considered in this work. Furthermore, $d\in\{6,7\}$ are reasonable compromises between accuracy and keeping the QUBO matrix as small as possible.
Figure 1: RMSE of the $d$ tabu search (red squares, $f_{\textup{tb}}^{\ast}$) and quantum annealing (green diamonds, $f_{\textup{qa}}^{\ast}$) solutions as a function of $d$. The RMSE of the exact, $d$-independent classical solution $f_{\textup{cl}}^{\ast}$ is shown as a gray dashed line. Lines between data points are shown to guide the eye. The underlying sample data set is the cubic one (Eq. 23), and triangular functions with $m=4$ are used for the approximation.
Last, we consider the case of varying $m$, i.e., how the solution quality depends on the number of triangular functions used for the curve fitting. This is an particularly interesting case because one usually chooses a sufficient numerical accuracy and simply increases $m$ to the point where the data is sufficiently good approximated. Curve fitting with triangular functions is well-suited for that because the result is a piecewise-linear approximation (a polygonal chain), which is often used to approximate more complex curves or if the functional form of the to-be-approximated data is not immediately obvious. In Fig. 2, we show the RMSE of the tabu (red squares), quantum annealing (green diamonds), and classical (gray circles) solutions as a function of $m$, where the sample data sets are the trigonometric (left diagram) and quadratic ones (right diagram); again, $n=64$, $d=8$, and $p=d-1=7$. First, we consider the left diagram. It shows the overall trend that the RMSE of the classical solution becomes smaller as $m$ increases. Initially, the tabu and quantum annealing solutions follow that trend. For $m\geq 10$ (which corresponds to a QUBO matrix of size $80\times 80$), however, the quantum annealing solution becomes worse with increasing $m$. The right diagram, though with a rather different underlying set of sample data (quadratic instead of trigonometric), shows the same qualitative behavior. Quantitatively, noticeable differences and a continuous worsening of the quantum annealing solutions with increasing $m$ already starts at $m=8$, which corresponds to a QUBO matrix size of $64\times 64$.
We also tested other polynomial sample data as well as varied the numerical coefficients of the sample data used for Fig. 2 and found qualitatively similar results. This indicates that a QUBO size of about $60\times 60$ to $80\times 80$ seems to be the range above which D-Wave’s current quantum annealing hardware fails to find a satisfying solution of the QUBO formulation of least squares curve fitting with triangular functions.
Figure 2: RMSE of the tabu search (red squares, $f_{\textup{tb}}^{\ast}$), quantum annealing (green diamonds, $f_{\textup{qa}}^{\ast}$) and classical (gray circles, $f_{\textup{cl}}^{\ast}$) solutions as a function of $m$. The underlying sample data are the trigonometric (left diagram) and quadratic (right diagram) ones, respectively.
Table 4: Optimal expansion coefficients of the individual Chebyshev polynomials of a quadratic fit, obtained classically via matrix inversion and by solving a QUBO via tabu search as well as quantum annealing. Also shown are the APE of the optimal coefficients obtained from the QUBO with respect to the classical ones.
$c_{0}^{\ast}$ $c_{1}^{\ast}$ $c_{2}^{\ast}$ APE $c_{0}^{\ast}$ [%] APE $c_{1}^{\ast}$ [%] APE $c_{2}^{\ast}$ [%]
Classical (matrix inversion) $0.544598$ $-0.031799$ $0.499257$
QUBO (tabu search) $0.523438$ $-0.039063$ $0.484375$ $3.89$ $22.84$ $2.98$
QUBO (quantum annealing) $0.445313$ $-0.101563$ $0.421875$ $18.23$ $219.39$ $15.50$
Figure 3: RMSE of the tabu search (red squares, $f_{\textup{tb}}^{\ast}$) and quantum annealing (green diamonds, $f_{\textup{qa}}^{\ast}$) solutions as a function of $d$. The RMSE of the exact, $d$-independent classical solution $f_{\textup{cl}}^{\ast}$ is shown as a gray dashed line. Lines between data points are shown to guide the eye. The underlying sample data set is the cubic one (Eq. 23), and Chebyshev polynomials with $m=4$ are used for the approximation.
@@ III.3 Chebyshev polynomials
Besides the triangular functions, we shall also consider Chebyshev polynomials of the first kind as defined in Eq. 3, $\phi_{j}=T_{j}$, i.e., the fit function is now given by
$f(x)=c_{0}+\sum_{j=1}^{m-1}c_{j}T_{j}(x)$ (30)
and is a polynomial of degree $m$.
In Tab. 4, we show our results for fitting a quadratic Chebyshev polynomial (i.e., $m=3$) to the quadratic sample data 22; again, $n=64$, $d=8$, and $p=d-1=7$, which results in a QUBO matrix of size $24\times 24$. While the tabu solution is in acceptable agreement with the classical one, the quantum annealing solution deviates significantly. We find the same result for the cubic data (also with varying numerical coefficients) with a cubic Chebyshev polynomial ($m=4$) as the fit function. In Fig. 3, we show the RMSE of the tabu (red squares) and quantum annealing (green diamonds) solutions $f_{\textup{tb}}^{\ast}$ and $f_{\textup{qa}}^{\ast}$, respectively, as a function of $d$, the number of digits that is used for the binary fixed-point format of the expansion coefficients. Furthermore, we show the RMSE of the classical solution, which is independent of $d$, as a gray dashed line. The underlying sample data is the cubic one and the Chebyshev fit is of degree three, i.e., $m=4$. We plot against $d$ because we would like to keep the degree of the Chebyshev polynomial fixed while letting the QUBO matrix increase in size. While the tabu solution shows the expected behavior (cf. Fig. 1), we are not able to obtain a satisfying quantum annealing solution for problem sizes larger than $20\times 20$. This is in stark contrast to our findings for piecewise-linear curve fitting using triangular functions.
@@
Figure 4: Heat map plots of exemplary QUBO matrices $Q$ of piecewise-linear approximation with triangular functions (left) and Chebyshev polynomial approximation (right). Here, we display the upper-triangularized versions of these matrices since every QUBO problem can be written in a way such that $Q$ is an upper-triangular matrix.
@@
This seems to hint toward the limitations of D-Wave’s current quantum annealing hardware for curve fitting with polynomials formulated as a QUBO. More precisely, the corresponding QUBO problem, which can be seen as a graph with logical qubits representing nodes that are connected by edges with weights given by the matrix elements of the QUBO matrix, needs to be mapped onto the so-called working graph of the quantum annealing device. Therefore, one might need several physical qubits to represent one logical qubit. Particularly for highly-connected (dense) QUBO problems, this could yield an Ising model that exhibits a complicated low-energy landscape where the ground state is difficult to access. This would explain why we were not successful in obtaining a feasible Chebyshev approximation as soon as its QUBO matrix is larger than the above mentioned size of about $20\times 20$. Such a matrix is, as apparent from the right diagram of Fig. 4, indeed relatively highly connected.
@@
We would like to emphasize that by solving the QUBO problems for piecewise-linear and Chebyshev approximations using a heuristic technique, e.g., tabu search or simulated annealing, we (almost) always find a solution in satisfying agreement with the classical one—even for sizes of the QUBO matrix for which the quantum annealing solution turned out to be unpractical. Thus, we are confident that a next-generation quantum annealing device will be capable of solving such QUBO problems.
Finally, we would like to mention that we always subjected the “raw” QUBO matrix to the quantum annealing process. It is well-known that dense QUBO problems are hard to solve, and a proper preprocessing of the QUBO matrix might improve the solution quality. Possible preprocessing steps are sparsification of the QUBO matrix or properly modifying/combining matrix elements [58, 59, 60], possibly resulting in a reduction in size and/or density of the problem. Another approach would be quantum annealing-inspired algorithms like SimCIM [61] or the NGQ solver [62], which seem to be more robust against dense QUBO problems [63, 64, 62]. However, these investigations are beyond the scope of this work and deferred to possible future work.